OpenStudy (anonymous):

find the equation of the tangent to the curve when x=2t^2+1 and y=(1/3)(t^3)-t and t=3

OpenStudy (anonymous):

tan( angel of curve) = y/x; so by t=3 we have y = 6 and x = 19 so tan(a o c ) = 6/19